What am I doing wrong? How would you set this up as an equation?
The Michigan plant of DynoTech Industries is four times as old as the Ohio plant. If the difference in the ages of the two plants is 9 years, what is the age of each?
x+4x=y-9
I didn't bring my text book, but I'll find out if I have another problem.
Thank You

Found 2 solutions by rchill, tutorcecilia:
Answer by rchill(405)

You've got a good start, but don't introduce a 2nd variable here. You've got x as the Michigan plant -- good; and 4x is the age of the Ohio plant -- very good. The difference in their ages is 9; that means that 4x-x=9. Solving for x, we get 3x=9, which simplifies to x=3. That means the Michigan plant is 3 years old and the Ohio plant is 4 times that, or 12 years old. Note the difference in ages is 9.

Answer by tutorcecilia(2152)

DynoTech Industries = 4 X Ohio plant = 4n
Ohio = n
If the difference in the ages of the two plants is 9 years, than
Dynotech - Ohio = 9
.
So,
(4 x Ohio) - (Ohio) = 9
4n - n = 9 [Solve for n]
3n=9
n=3 is the age of the Ohio plant
.
checking:
Dynatech - Ohio = 9
(4)(3)-(3) = 9
12-3=9
9=9 [Checks out]